# Dissenting Explanations: Leveraging Disagreement to Reduce Model Overreliance

### v2

@@ Introduction
The development of increasingly capable AI systems has motivated the adoption of AI-assisted decision-making. In high-stakes settings such as loan approval and patient diagnosis, it is imperative for humans to understand how any given model came to its decision. However, with the success of deep learning, many large state-of-the-art models are not easily interpretable. Thus, explainability (XAI) methods are crucial for providing justification for the decisions of black-box models. Such explanations justify a model’s prediction on a singular input example, and their goal is to provide accurate information while also being succinct and easy for humans to parse (Burkart and Huber 2020).
@@
Wang and Yin (2021) list the desiderata of explanations as (1) model understanding, (2) helping people recognize model uncertainty, and (3) calibrating trust for AI models. Towards the goal of understanding a single model, recent works have shown that explanations generated from different methods based on the same instance can conflict (Han, Srinivas, and Lakkaraju 2022; Krishna et al. 2022). While multiple explanations for the same model question the validity of explainability tools, diverse and even conflicting explanations for the same decision might better calibrate trust and illustrate uncertainty in a model-aided decision-making setting. Instead of rejecting explanations altogether, the existence of multiple plausible explanations motivates the perspective that explanations can be treated as arguments supporting a given model recommendation in the decision-making process.
@@
With the framework of explanations as arguments, we may naturally construct a courtroom analogy, in which human decision makers are the judges deciding whether the model prediction is trustworthy. When a singular explanation is provided, a decision-maker may be unduly influenced to trust the prediction. Indeed, Bansal et al. (2021) show that when explanations are provided, humans are more likely to follow a model decision regardless of whether the model is correct. Thus, while an explanation provides a supporting argument for a prediction, we must also provide alternative arguments, arguing against the model prediction, in order to accommodate meticulous human decision-making. In the context of a consequential legal decision, presenting both sides amounts to procedural due process.
@@
In this paper, we introduce the notion of dissenting explanations: explanations for an opposing model prediction to some reference model. To illustrate the importance of these explanations, we focus, from human study to proposed techniques, on a single task: deceptive hotel reviews classification. We perform a study to show that, on this difficult-to-verify task, dissenting explanations indeed reduce model overreliance without reducing the accuracy of the human predictions. Finally, since dissenting explanations are a useful tool for reducing overreliance, even outside the context of existing model multiplicity, we develop methods to induce predictive multiplicity and create dissenting explanations. We present techniques for generating global disagreement with respect to any black-box model, as well as local disagreement on any instance; these methods achieve disagreement without sacrificing model accuracy  [^1] Full paper with appendices available at https://arxiv.org/abs/2307.07636.
Related Work
@@
Post-hoc explanations can be elicited from black box models through techniques including perturbation-based methods (Ribeiro, Singh, and Guestrin 2016; Lundberg and Lee 2017) and gradient-based methods (Selvaraju et al. 2017; Smilkov et al. 2017). However, when applying such techniques to the same example, inconsistent and conflicting explanations for feature importance may arise. Surveying data scientists, Krishna et al. (2022) found disagreements in feature explanations and find that more complex models exhibit higher disagreement.
@@
For trustworthy predictions, humans may expect similar predictions from similarly accurate models. Yet models with similar accuracy still exhibit different predictions (Marx, Calmon, and Ustun 2020). Brunet, Anderson, and Zemel (2022) show that similar accuracy models can yield vastly different explanations due to different random seeds and hyperparameters. Other works in robustness purposefully generate disagreement between models (Pang et al. 2019; Rame and Cord 2021).
@@
Among tasks where neither humans nor AI routinely achieves perfect performance, Lai and Tan (2019) use AI predictions and explanations to help human participants with detecting deceptive hotel reviews and find that human performance was improved with AI predictions with explanations. Vasconcelos et al. (2022) also find that explanations actually reduce overreliance in their set of maze task experiments. In contrast, Bansal et al. (2021) study common sense tasks including review sentiment classification and found that explanations increased accuracy when the AI model was correct but decreased accuracy when the AI model was wrong. An adjacent line of work studies AI debate: two large language models are prompted to recursively try to convince a human judge to take their side (Michael et al. 2023).
We investigate the effect of also showing the explanation of a dissenting model in reducing overreliance in settings where AI surpasses human performance, but human decision-makers may need to make the final decision (Lai and Tan 2019; Lundberg and Lee 2017). In these settings, the goal is to provide AI predictions with explanations to humans as a tool rather than removing humans from decision-making altogether. The closest prior work to ours includes a position paper arguing for decision support systems to provide support for and against decisions (Miller 2023) and the effect of using different explanations from the same model (Bansal et al. 2021). In contrast, we examine differing independent predictions and accompanying explanations from different models in the Rashomon set (Fisher, Rudin, and Dominici 2019) with the goal of improving human decision-making.
Model and Framework
@@
Definition 1 (Dissenting Explanation).
@@
Definition 2 (Global Predictive Disagreement).
@@
While disagreement can be maximized by models that disagree on every example, we care about the setting of models with similarly high accuracy, also described as the Rashomon set (Fisher, Rudin, and Dominici 2019).
Following prior work studying the overreliance of humans on AI predictions (Vasconcelos et al. 2022), we define overreliance as how much human decisions mirror AI suggestions when the AI is incorrect: ${\mathbb{E}[h(x)=f(x)|f(x)\neq y]}$ where $h$ represents the human decision.
@@
Motivating Study: The Importance of Dissenting Explanations
@@
Figure 1: The process of generating explanations for the 4 conditions. The review text is for conceptual understanding only.
@@
We focus on the setting of assistive AI: the setting where AI on average might perform better than humans but it is critical for humans to be the final decision maker. This is different from prior works, which focused on tasks either with “verifiable” answers given the explanation (Vasconcelos et al. 2022) or tasks where humans and AI perform approximately equally in order to measure collaboration potential (Bansal et al. 2021). We use the term “verifiable” to describe tasks where explanations can verify that a model prediction is true (e.g. a maze task explanation provides a path to complete the maze or a math task explanation provides the step-by-step solution). We specifically consider explanations that are not verifiable proofs of the correct label but rather arguments for the model predictions, as these are the standard explanation forms available for complex model predictions (Burkart and Huber 2020). Since the effects of explanations differ widely from task to task (Wang and Yin 2021), we focus on finding a specific, suitable task for studying dissenting explanations. We set the following requirements for our target task [^2] We elaborate on these desiderata and discuss the candidate tasks we considered in the supplementary materials.:
@@
Based on our requirements, we decided to use the Chicago Deceptive Reviews dataset (Ott et al. 2011): a dataset of 1600 one-paragraph reviews of hotels in Chicago, where half the reviews are genuine reviews from TripAdvisor, and the other half were written by crowd workers that have only seen the name and website of the hotel. The goal of the task is to distinguish between real and deceptive reviews; a prior study found that humans on their own get at most 62% accuracy on this task, while a linear SVM achieved around 87% accuracy (Lai and Tan 2019). Furthermore, there exists a ground truth label: whether a review is deceptive or real. The explanations were in the form of highlighting the words selected by the feature attribution explainer; these words serve as an argument to the participant, convincing them to select a certain label without giving a complete proof of the correct answer.
To test our hypotheses, we design a study in which human participants attempt to categorize these hotel reviews. Participants are presented with 20 hotel reviews, each of which is real or deceptive, and are instructed to decide which reviews are real. They are assisted by AI predictions or explanations, where the existence or type of explanation varies based on the condition participants are assigned to. Participants are warned in the beginning that the AI predictions are not always correct. Based on response quality in pilot studies, participants are also given a set of heuristics for identifying deceptive reviews, developed by prior work on this task (Lai, Liu, and Tan 2020). We also survey users, post-task, about task difficulty, the helpfulness of AI suggestions, and their trust in the AI suggestions. Finally, we included an optional open-ended question about how the AI suggestions helped them complete the task.
@@
To properly benchmark against prior work (Lai and Tan 2019), we use the same linear SVM trained on TF-IDF of unigrams with English stop words removed as our reference model $f$ with 87% accuracy. We also train an alternative 3-layer neural network model based on the exact same pre-processing which achieves 79% accuracy as the alternative model $g$. We use LIME (Ribeiro, Singh, and Guestrin 2016) to generate local explanations for each model using the Top-k features (k=15 was chosen based on the number of unique tokens per example for this dataset).
We used LIME for generating explanations in order to (1) remain consistent against prior work XAI studies in the text domain (Bansal et al. 2021), and (2) benchmark against (Lai and Tan 2019) who use linear model features (i.e. we compared our LIME features to the weights of the linear SVM model and found meaningful overlap [^3] See details in in the supplementary materials). To find dissenting explanations, we used examples in the test set where the neural network model disagreed with the linear SVM model. The two models disagreed on $10\%$ of examples (32 examples). We sub-sampled these examples for an even balance of examples that the linear SVM (the reference model) predicted correctly and incorrectly. This selection process yields a subset of examples that result in $\sim 60\%$ accuracy; both for the reference model $f$ and human baseline.
@@ Results
For each participant in the study, we measured their accuracy as the fraction of reviews they categorized correctly, out of the 20 total reviews. We measured overreliance as the fraction of reviews they agreed with the model prediction on, out of the 8 reviews the model predicted incorrectly. These results, averaged over users of each of the four conditions, are displayed in Figure 1(a) and Figure 1(b). Since our task involves binary labels, we account for random agreement by also measuring Cohen’s $\kappa$ between a participant and the model’s predictions in Figure 1(c) (McHugh 2012).
@@
Finding Dissenting Explanations
Motivated by the utility of dissenting explanations, we present methods for producing disagreement in models. While many techniques may be developed to create model multiplicity, we focus on simple objective-based approaches that are effective for our task of interest similar to techniques from robustness and diverse ensembles (Pang et al. 2019; Rame and Cord 2021) [^5] Additional experiments for these techniques on other datasets are included in the supplementary materials. While prior works have focused on predictive multiplicity, a clear mapping between predictive and explanation multiplicity has not been presented. In this section, we present and compare methods for increasing predictive multiplicity through the lens of explanations.
Global Disagreement: A Model Agnostic Approach
@@
Problem 4.
Global Disagreement Given reference model $f$ and training data $D$, find some $g$ such that $\delta_{D}(f,g)$ (Definition 2) is maximized while $\operatorname{Err}_{D_{\text{test}}}(f)\approx\operatorname{Err}_{D_{\text{test}}}(g)$.
@@
For comparing dissenting explanations to the original explanation, we use a similar set of metrics as explanation agreement (Krishna et al. 2022). We consider three agreement metrics TopK, TopKPos, and TopKNeg:
@@
Local Disagreement: Generating a Dissenting Explanation for Any Input
@@
Here we know the exact test instance for which we want a different prediction. We both consider the scenario of additional access to only the training set and the scenario of additional access to only the trained model. For the former, we use a linear SVM to demonstrate the efficacy of flipping the test instance label by reducing the training size $|D|$. As the training data size decreases, the influence of the test instance increases, improving the success rate of generating a different prediction for that instance. Table 1(a) shows that Problem 6 can be solved with high probability (i.e. $>90\%$) without sacrificing significant model accuracy for the deceptive reviews task. For the scenario where we only have additional access to the trained model (i.e. a neural network fitted to the training set), we retrain directly to minimize the loss on the test instance and measure how many iterations are required to change the label. Table 1(b) describes the distribution over iterations required to flip an example label. This method is also effective in flipping the label for roughly $80\%$ of the test set examples while still maintaining $\sim 88\%$ accuracy.
Discussion
In this work, we take a holistic approach by first motivating the need for dissenting explanations through a human study to measure overreliance. For our deceptive reviews task, a task with a ground truth label but no method for direct verification, we demonstrate the utility of dissenting explanations in reducing overreliance. Our results complement existing work on the benefits of explanations (Vasconcelos et al. 2022) by exploring more ambiguous tasks.
@@ Discussion
Our work, on a single task, cannot make claims about the effectiveness of dissenting explanations in general. The subset of examples we examine is difficult for the reference model, thus providing a single explanation does not significantly assist human decision-making; future work should use a stronger reference model with more examples. A promising direction of future work is to explore what tasks dissenting explanations best aid and other types of dissenting explanations involving counterfactual explanations.
@@ Appendix A Additional Experiments
In our main text, we focus on the Hotel reviews dataset we use for our study. To demonstrate the effectiveness of our proposed approaches, we run the same set of experiments both for global and local disagreement on the UCI Adult dataset from Folktables (Ding et al. 2021). We use the Income task for Hawaii from 2018.
@@ Appendix B Human Experiment Details
In order to properly analyze how dissenting-explanations affect human overreliance, the task we use must allow humans to learn from AI explanations, without explanations completely solving the problem for the human participants. Prior works focused on tasks either with verifiable answers given the explanation (Vasconcelos et al. 2022) or tasks where humans and AI perform approximately equally in order to measure collaboration potential (Bansal et al. 2021). We focus on assistive AI: the setting where AI on average might perform better than humans but it is critical for humans to be the final decision maker. Furthermore, we specifically consider explanations that cannot fully verify the model prediction but rather makes an argument for the model’s prediction. Thus, we had the following criteria in selecting a task:
@@
With these criteria in mind, we review existing tasks where the value of explanations has been tested. Table 5 summarizes the explainability tasks that have been studied within the context of AI decision-making as well as the reasons that these tasks did not match our criteria. The only prior work presenting a task that matched our criteria is by Buçinca et al. (2020). They provide other examples in as inductive support for an AI’s decision as well as ingredient lists for deductive support. For the purpose of our study, we focus on deductive support from explanations with are generated by real-world explainability tools. It is unclear what model disagreement would look like on this task so we leave the study of dissenting explanations on this task to future work.
Thus we end up with only deceptive reviews as the task with precedent in prior work which fits all of our criteria. Moreover, the human-only baseline on this task has already been tested to be 51.1% (Lai and Tan 2019).
Task
Reason for Elimination
Amazon Reviews (Bansal et al. 2021)
Machine accuracy worse than humans (Reason 1)
LSAT (Bansal et al. 2021)
Machine accuracy worse than humans (Reason 1)
Income Prediction (Zhang, Liao, and Bellamy 2020)
No objectively correct label (Reason 3)
Loan Defaults (Green and Chen 2019)
No objectively correct label (Reason 3)
Hypoxemia risk (Lundberg et al. 2018)
Explanation understanding requires domain expertise (Reason 4)
Maze (Vasconcelos et al. 2022)
No significant model disagreement and explanation provides proof/verification for prediction (Reason 2 & 4)
Deceptive Reviews (Lai and Tan 2019)
N/A (Chosen Task)
@@
Each participant was presented the same 20 reviews and AI predictions, and asked to categorize each review as deceptive or real. The only difference between the questions was in the type of AI explanation presented to the participants. We provide an example question from the survey, under each of the 4 conditions. The question and explanation for $C_{0}$ can be seen in 5(a), the question and explanation for $C_{1}$ can be seen in 5(b), and the prompts for $C_{2}$ and $C_{3}$ are in 6(a) and 6(b), respectively.
@@
Among the many explanation methods, we use LIME to best adhere to the closed prior work to our endeavor (Bansal et al. 2021). However, since this work does not specifically look at the deceptive reviews task, we run a few validation techniques to ensure that the explanations are meaningful. Note we run this validation on just the base model since lower-quality base model explanations may reduce trust and result in an over-report of the effect in H2 (reduction in overreliance) that we are trying to measure.
@@
Lai and Tan (2019) use the actual weights of a linear model for their feature explanations for their human study. To validate that LIME features are sufficiently close to these features, we compute overlap metrics between lime features and our linear SVM features. We observe top feature overlap TopK of 56% and predictive feature overlap of 60%. This indicates that there is a significant overlap in top features between Linear SVM and LIME (Compared to results in (Krishna et al. 2022)).
Human Baseline
After running this study, we also ran a human baseline, with 45 participants answering the questions with no model assistance. This baseline was run in exactly the same format as the previous study, except with no model assistance, and correspondingly, the questions related to trust in the model and the instructions related to model assistance removed. The participants in the baseline had an accuracy of $0.61\pm 0.017$, and a Cohen’s Kappa agreement score of $0.115\pm 0.031$ with the model. The accuracy aligns with the prior work (Lai and Tan 2019), which found that humans on their own get an accuracy of around $62\%$ on the Chicago dataset as a whole. Moreover, the Cohen’s Kappa agreement score is also expected, as we would expect to have low agreement with a model that the respondents do not see, but a nonzero agreement due to the fact that humans and the model may pick up on similar signs of deceptive reviews.
The overall accuracy does not significantly change between the human-only condition and any of $C0$ through $C4$ (we get $p$-values between 0.15 and 0.45 when comparing accuracy to each of these conditions). This is perhaps expected, due to the fact that the reference model also gets a $60\%$ accuracy on these questions. However, the answers to these questions do significantly change between conditions; the model agreement for participants in $C1$ (who were presented with one explanation) increases to $0.34\pm 0.08$, this gives us a $p$-value of $51.1e-$ 5 when compared with the human-only baseline. In contrast, $C2$ (participants who were presented with dissenting explanations), show a similar agreement to the human-only baseline, and a question-level analysis shows that $C2$ only significantly differs from the baseline on one of the questions; this is expected by the definition of $p$-value for a set of 20 questions where all the answers are drawn from the same distribution.
Thus, these results indicate that humans have a reasonable accuracy on this subset of questions on their own, and though this accuracy does not change, they tend to agree with the model (on both correct and incorrect answers) in conditions where they are presented no or one explanation, and learn to not overrely on the model and follow their own intuition more when presented with dissenting explanations. This is further evidenced by some of the comments from participants in $C2$, such as “I didn’t feel like most of the keywords the AI was highlighting were really ones that helped me determine real or fake. So I would consult the AI briefly, but then go with my own determination more based on the details given.”
Based on this baseline, an interesting direction for future study would be to run this experiment on a different task, and with a different form of explanation, so that the model-only accuracy is higher than the human-only accuracy, even on the subset of questions given, and so that the explanations are more human-readable, allowing for further granularity in the difference between the four conditions.

### v3

@@ 1 Introduction
The development of increasingly capable AI systems has motivated the adoption of AI-assisted decision-making. In high-stakes settings such as loan approval and patient diagnosis, it is imperative for humans to understand how any given model came to its decision. However, with the success of deep learning, many large state-of-the-art models are not easily interpretable. Thus, explainability (XAI) methods are crucial for providing justification for the decisions of black-box models. Such explanations justify a model’s prediction on a singular input example, and their goal is to provide accurate information while also being succinct and easy for humans to parse (burkart2020survey).
@@
wang2021explanations list the desiderata of explanations as (1) model understanding, (2) helping people recognize model uncertainty, and (3) calibrating trust for AI models. Towards the goal of understanding a single model, recent works have shown that explanations generated from different methods based on the same instance can conflict (han2022explanation; krishna2022disagreement). While multiple explanations for the same model question the validity of explainability tools, diverse and even conflicting explanations for the same decision might better calibrate trust and illustrate uncertainty in a model-aided decision-making setting. Instead of rejecting explanations altogether, the existence of multiple plausible explanations motivates the perspective that explanations can be treated as arguments supporting a given model recommendation in the decision-making process.
@@
With the framework of explanations as arguments, we may naturally construct a courtroom analogy, in which human decision makers are the judges deciding whether the model prediction is trustworthy. When a singular explanation is provided, a decision-maker may be unduly influenced to trust the prediction. Indeed, bansal2021does show that when explanations are provided, humans are more likely to follow a model decision regardless of whether the model is correct. Thus, while an explanation provides a supporting argument for a prediction, we must also provide alternative arguments, arguing against the model prediction, in order to accommodate meticulous human decision-making. In the context of a consequential legal decision, presenting both sides amounts to procedural due process.
@@
In this paper, we introduce the notion of dissenting explanations: explanations for an opposing model prediction to some reference model. To illustrate the importance of these explanations, we focus, from human study to proposed techniques, on a single task: deceptive hotel reviews classification. We perform a study to show that, on this difficult-to-verify task, dissenting explanations indeed reduce model overreliance without reducing the accuracy of the human predictions. Finally, since dissenting explanations are a useful tool for reducing overreliance, even outside the context of existing model multiplicity, we develop methods to induce predictive multiplicity and create dissenting explanations. We present techniques for generating global disagreement with respect to any black-box model, as well as local disagreement on any instance; these methods achieve disagreement without sacrificing model accuracy.
2 Related work
@@
Post-hoc explanations can be elicited from black box models through techniques including perturbation-based methods (ribeiro2016should; lundberg2017unified) and gradient-based methods (selvaraju2017grad; smilkov2017smoothgrad). However, when applying such techniques to the same example, inconsistent and conflicting explanations for feature importance may arise. Surveying data scientists, krishna2022disagreement found disagreements in feature explanations and find that more complex models exhibit higher disagreement.
@@
For trustworthy predictions, humans may expect similar predictions from similarly accurate models. Yet models with similar accuracy still exhibit different predictions (marx2020predictive). brunetimplications show that similar accuracy models can yield vastly different explanations due to different random seeds and hyperparameters. Other works in robustness purposefully generate disagreement between models (pang2019improving; rame2021dice).
@@
Among tasks where neither humans nor AI routinely achieves perfect performance, lai2019human use AI predictions and explanations to help human participants with detecting deceptive hotel reviews and find that human performance was improved with AI predictions with explanations. vasconcelos2022explanations also find that explanations actually reduce overreliance in their set of maze task experiments. In contrast, bansal2021does study common sense tasks including review sentiment classification and found that explanations increased accuracy when the AI model was correct but decreased accuracy when the AI model was wrong.
We investigate the effect of also showing the explanation of a dissenting model in reducing overreliance in settings where AI surpasses human performance, but human decision-makers may need to make the final decision (lai2019human; lundberg2017unified). In these settings, the goal is to provide AI predictions with explanations to humans as a tool rather than removing humans from decision-making altogether. Our setting differs from (bansal2021does) in that we examine differing independent predictions and accompanying explanations from different models in the Rashomon set (fisher2019all) with the goal of improving human decision-making.
3 Model and framework
@@
Definition 3.1 (Dissenting Explanation).
@@
Definition 3.2 (Global predictive disagreement).
@@
While disagreement can be maximized by models that disagree on every example, we care about the setting of models with similarly high accuracy, also described as the Rashomon set (fisher2019all).
Following prior work studying the overreliance of humans on AI predictions (vasconcelos2022explanations), we define overreliance as how much human decisions mirror AI suggestions when the AI is incorrect: ${\mathbb{E}[h(x)=f(x)|f(x)\neq y]}$ where $h$ represents the human decision.
@@
4 Motivating study: the importance of dissenting explanations
@@
Figure 1: The process of generating explanations for the 4 conditions.
@@
We focus on the setting of assistive AI: the setting where AI on average might perform better than humans but it is critical for humans to be the final decision maker. This is different from prior works, which focused on tasks either with “verifiable” answers given the explanation (vasconcelos2022explanations) or tasks where humans and AI perform approximately equally in order to measure collaboration potential (bansal2021does). We use the term “verifiable” to describe tasks where explanations can verify that a model prediction is true (e.g. a maze task explanation provides a path to complete the maze or a math task explanation provides the step-by-step solution). We specifically consider explanations that are not verifiable proofs of the correct label but rather arguments for the model predictions, as these are the standard explanation forms available for complex model predictions (burkart2020survey). Since the effects of explanations differ widely from task to task (wang2021explanations), we focus on finding a specific, suitable task for studying dissenting explanations. We set the following requirements for our target task [^1] We elaborate on these desiderata and discuss the candidate tasks we considered in the supplementary materials.:
@@
Based on our requirements, we decided to use the Chicago Deceptive Reviews dataset (ott2011finding): a dataset of 1600 one-paragraph reviews of hotels in Chicago, where half the reviews are genuine reviews from TripAdvisor, and the other half were written by crowd workers that have only seen the name and website of the hotel. The goal of the task is to distinguish between real and deceptive reviews; a prior study found that humans on their own get at most 62% accuracy on this task, while a linear SVM achieved around 87% accuracy (lai2019human). Furthermore, there exists a ground truth label: whether a review is deceptive or real. The explanations were in the form of highlighting the words selected by the feature attribution explainer; these words serve as an argument to the participant, convincing them to select a certain label without giving a complete proof of the correct answer.
To test our hypotheses, we design a study in which human participants attempt to categorize these hotel reviews. Participants are presented with 20 hotel reviews, each of which is real or deceptive, and are instructed to decide which reviews are real. They are assisted by AI predictions or explanations, where the existence or type of explanation varies based on the condition participants are assigned to. Participants are warned in the beginning that the AI predictions are not always correct. Based on response quality in pilot studies, participants are also given a set of heuristics for identifying deceptive reviews, developed by prior work on this task (lai2020chicago). We also survey users, post-task, about task difficulty, the helpfulness of AI suggestions, and their trust in the AI suggestions. Finally, we included an optional open-ended question about how the AI suggestions helped them complete the task.
@@
To properly benchmark against prior work (lai2019human), we use the same linear SVM trained on TF-IDF of unigrams with English stop words removed as our reference model $f$ with 87% accuracy. We also train an alternative 3-layer neural network model based on the exact same pre-processing which achieves 79% accuracy as the alternative model $g$. We use LIME (ribeiro2016should) to generate local explanations for each model using the Top-k features (k=15 was chosen based on the number of unique tokens per example for this dataset).
We used LIME for generating explanations in order to (1) remain consistent against prior work XAI studies in the text domain (bansal2021does), and (2) benchmark against (lai2019human) who use linear model features (i.e. we compared our LIME features to the weights of the linear SVM model and found meaningful overlap [^2] See details in in the supplementary materials). In order to find dissenting explanations, we used examples in the test set where the neural network model disagreed with the linear SVM model. The two models disagreed on $10\%$ of examples (32 examples). We sub-sampled these examples for an even balance of examples that the linear SVM (the reference model) predicted correctly and incorrectly.
@@ Results
For each participant in the study, we measured their accuracy as the fraction of reviews they categorized correctly, out of the 20 total reviews. We measured overreliance as the fraction of reviews they agreed with the model prediction on, out of the 8 reviews the model predicted incorrectly. These results, averaged over users of each of the four conditions, are displayed in Figure 2(a) and Figure 2(b). Since our task involves binary labels, we account for random agreement by also measuring Cohen’s $\kappa$ between a participant and the model’s predictions in Figure 2(c) (mchugh2012interrater).
@@
5 Finding dissenting explanations
Motivated by the utility of dissenting explanations, we present methods for producing disagreement in models. While many techniques may be developed to create model multiplicity, we focus on simple objective-based approaches that are effective for our task of interest similar to techniques from robustness and diverse ensembles (pang2019improving; rame2021dice) [^4] Additional experiments for these techniques on other datasets are included in the supplementary materials. While prior works have focused on predictive multiplicity, a clear mapping between predictive and explanation multiplicity has not been presented. In this section, we present and compare methods for increasing predictive multiplicity through the lens of explanations.
Global model disagreement: a model agnostic approach
@@ 5 Finding dissenting explanations
Problem 5.1.
Global Disagreement Given reference model $f$ and training data $D$, find some $g$ such that $\delta_{D}(f,g)$ (Definition 3.2) is maximized while $\operatorname{Err}_{D_{\text{test}}}(f)\approx\operatorname{Err}_{D_{\text{test}}}(g)$.
@@
For comparing dissenting explanations to the original explanation, we use a similar set of metrics as explanation agreement (krishna2022disagreement). We consider three agreement metrics TopK, TopKPos, and TopKNeg:
@@
Here we know the exact test instance for which we want a different prediction. We both consider the scenario of additional access to only the training set and the scenario of additional access to only the trained model. For the former, we use a linear SVM to demonstrate the efficacy of flipping the test instance label by reducing the training size $|D|$. As the training data size decreases, the influence of the test instance increases, improving the success rate of generating a different prediction for that instance. Table 2(a) shows that Problem 5.3 can be solved with high probability (i.e. $>90\%$) without sacrificing significant model accuracy for the deceptive reviews task. For the scenario where we only have additional access to the trained model (i.e. a neural network fitted to the training set), we retrain directly to minimize the loss on the test instance and measure how many iterations are required to change the label. Table 2(b) describes the distribution over iterations required to flip an example label. This method is also effective in flipping the label for roughly $80\%$ of the test set examples while still maintaining $\sim 88\%$ accuracy.
6 Discussion
In this work, we take a holistic approach by first motivating the need for dissenting explanations through a human study to measure overreliance. For our deceptive reviews task, a task with a ground truth label but no method for direct verification, we demonstrate the utility of dissenting explanations in reducing overreliance. Our results complement existing work on the benefits of explanations (vasconcelos2022explanations) by exploring more ambiguous tasks.
@@ 6 Discussion
We focus on a single task in our human study and thus we cannot make claims about the effectiveness of dissenting explanations more generally. A promising direction of future work is to explore what tasks dissenting explanations best aid and other types of dissenting explanations involving counterfactual explanations. Furthermore, while we presented a preliminary intuition on explanation disagreement from one vs two models, more studies are required to understand whether these results are due to a difference in the selected features between the two conditions, or simply a difference in framing.
@@ Appendix A Additional Experiments
In our main text, we focus on the Hotel reviews dataset we use for our study. To demonstrate the effectiveness of our proposed approaches, we run the same set of experiments both for global and local disagreement on the UCI Adult dataset from Folktables (ding2021retiring). We use the Income task for Hawaii from 2018.
@@ Appendix B Human Experiment Details
In order to properly analyze how dissenting-explanations affect human overreliance, the task we use must allow humans to learn from AI explanations, without explanations completely solving the problem for the human participants. Prior works focused on tasks either with verifiable answers given the explanation (vasconcelos2022explanations) or tasks where humans and AI perform approximately equally in order to measure collaboration potential (bansal2021does). We focus on assistive AI: the setting where AI on average might perform better than humans but it is critical for humans to be the final decision maker. Furthermore, we specifically consider explanations that cannot fully verify the model prediction but rather makes an argument for the model’s prediction. Thus, we had the following criteria in selecting a task:
@@
With these criteria in mind, we review existing tasks where the value of explanations has been tested. Table 5 summarizes the explainability tasks that have been studied within the context of AI decision-making as well as the reasons that these tasks did not match our criteria. The only prior work presenting a task that matched our criteria is by buccinca2020proxy. They provide other examples in as inductive support for an AI’s decision as well as ingredient lists for deductive support. For the purpose of our study, we focus on deductive support from explanations with are generated by real-world explainability tools. It is unclear what model disagreement would look like on this task so we leave the study of dissenting explanations on this task to future work.
Thus we end up with only deceptive reviews as the task with precedent in prior work which fits all of our criteria. Moreover, the human-only baseline on this task has already been tested to be 51.1% (lai2019human).
Task Reason for Elimination
Amazon Reviews (bansal2021does) Machine accuracy worse than humans (Reason 1)
LSAT (bansal2021does) Machine accuracy worse than humans (Reason 1)
Income Prediction (zhang2020effect) No objectively correct label (Reason 3)
Loan Defaults (green2019principles) No objectively correct label (Reason 3)
Hypoxemia risk (lundberg2018explainable) Explanation understanding requires domain expertise (Reason 4)
Maze (vasconcelos2022explanations) No significant model disagreement and explanation provides proof/verification for prediction (Reason 2 & 4)
Deceptive Reviews (lai2019human) N/A (Chosen Task)
@@
Each participant was presented the same 20 reviews and AI predictions, and asked to categorize each review as deceptive or real. The only difference between the questions was in the type of AI explanation presented to the participants. We provide an example question from the survey, under each of the 4 conditions. The question and explanation for $C_{0}$ can be seen in 6(a), the question and explanation for $C_{1}$ can be seen in 6(b), and the prompts for $C_{2}$ and $C_{3}$ are in 7(a) and 7(b), respectively.
@@
Among the many explanation methods, we use LIME to best adhere to the closed prior work to our endeavor (bansal2021does). However, since this work does not specifically look at the deceptive reviews task, we run a few validation techniques to ensure that the explanations are meaningful. Note we run this validation on just the base model since lower-quality base model explanations may reduce trust and result in an over-report of the effect in H2 (reduction in overreliance) that we are trying to measure.
@@
lai2019human use the actual weights of a linear model for their feature explanations for their human study. To validate that LIME features are sufficiently close to these features, we compute overlap metrics between lime features and our linear SVM features. We observe top feature overlap TopK of 56% and predictive feature overlap of 60%. This indicates that there is a significant overlap in top features between Linear SVM and LIME (Compared to results in (krishna2022disagreement)).